# Multi-modal Misinformation Detection: Approaches, Challenges and Opportunities

### v6

@@ Abstract.
As social media platforms are evolving from text-based forums into multi-modal environments, the nature of misinformation in social media is also transforming accordingly. Taking advantage of the fact that visual modalities such as images and videos are more favorable and attractive to the users and textual contents are sometimes skimmed carelessly, misinformation spreaders have recently targeted contextual connections between the modalities e.g., text and image. Hence many researchers have developed automatic techniques for detecting possible cross-modal discordance in web-based content. We analyze, categorize and identify existing approaches in addition to challenges and shortcomings they face in order to unearth new research opportunities in the field of multi-modal misinformation detection.
@@ 1. Introduction
Nowadays billions of multi-modal posts containing texts, images, videos, soundtracks etc. are shared throughout the web, mainly via social media platforms such as Facebook, Twitter, Snapchat, Reddit, Instagram, YouTube and so on and so forth. While the combination of modalities allows for more expressive, detailed and user-friendly content, it brings about new challenges, as it is harder to accommodate uni-modal solutions to multi-modal environments.
@@
Similar to other multi-modal tasks, it is harder and more challenging to detect fake news on multi-modal platforms, as it requires not only the evaluation of each modality, but also cross-modal connections and credibility of the combination as well. This becomes even more challenging when each modality e.g., text or image is credible but the combination creates misinformative content. For instance, a COVID-19 anti-vaccination misinformation [^1] Misinformation” is false information that spreads unintentionally, whereas the term ”Disinformation” refers to false information that malicious users share intentionally and often strategically to affect other audiences’ behaviors toward social, political, and economic events. In this work, regardless of spreaders’ intention, we refer to all sorts of false news i.e., misinformation and disinformation as ”Misinformation” or ”Fake News” interchangeably. post can have text that reads ”vaccines do this” and then attaches a graphic image of a dead person. In this case, although image and text are not individually misinformative, taken together create misinformation.
@@
Over the past decade, several detection models (Shu et al., 2017, 2020b; Islam et al., 2020; Cai et al., 2020) have been developed to detect misinformation. However, the majority of them leverage only a single modality for misinformation detection, e.g., text (Horne and Adali, 2017; Wu et al., 2017; Guacho et al., 2018; Shu et al., 2020a) or image (Huh et al., 2018; Qi et al., 2019; Abdali et al., 2021a; Choudhary and Arora, 2021), which miss the important information conveyed by other modalities. There are existing works (K. Shu and Liu, 2019; Shu et al., 2019a; Abdali et al., 2021b, 2020; Hakak et al., 2021) that leverage ensemble methods which create multiple models for each modality and then combine them to produce improved results. However, in many multi-modal misinformative content, individual modalities combined loosely together are insufficient to identify fake news and as a result, the joint model also fails.
@@
Nevertheless, in recent years, machine learning scientists have come up with different techniques for cross-modal fake news detection which combine information from multiple modalities leveraging cross-modal information, such as the consistency and meaningful relationships between different modalities. Study and analysis of these techniques and identifying existing challenges will give a clearer picture of the state of knowledge on multi-modal misinformation detection and open the door to new opportunities in this field.
@@
Even though there are a number of valuable surveys on fake news detection (Shu et al., 2017; Kumar and Shah, 2018; da Silva et al., 2019), very few of them focus on multi-modal techniques (Alam et al., 2021; Alonso-Bartolome and Segura-Bedmar, 2021). Since the number of proposed techniques on multi-modal fake news detection has been increasing immensely, the necessity of a comprehensive survey on existing techniques, datasets and emerging challenges is felt more than ever. With that said, in this work, we aim to conduct a comprehensive study on fake news detection in multi-modal environments.
@@
To this end, we classify multi-modal misinformation detection study into three different directions as follows:
@@
  • Multi-modal Data Study: In this direction, the goal is to collect multi-modal fake news data e.g., image, text, social context etc. from different sources of information and use fact-checking resources to evaluate the veracity of collected data and annotate them accordingly. Comparison and analysis of existing datasets as well as bench-marking are other tasks that are under the umbrella of this category.
@@
  • Multi-modal Feature Study: In this direction, the main objective is to identify meaningful connections between different data modalities which are often manipulated by misinformation spreaders to falsify, imposter or exaggerate original information. These meaningful connections may be used as clues for detecting misinformation in multi-modal environments such as social media posts. Another goal of this direction is to study and develop strategies for fusing feature of different modalities and create information-rich multi-modal feature vectors.
@@
  • Multi-modal Model Study: In this direction, the main focus is on the development of efficient multi-modal machine learning solutions to detect misinformation leveraging multi-modal features and clues. Proposing new techniques and approaches, in addition to improving performance, scalability, interpretability and explicability of machine learning models are some of the common tasks in this direction.
@@
These three directions create a pipeline in multi-modal misinformation study i.e., output of each study provides an input for the next one. A summary of the aforementioned directions is illustrated in Fig. 1. In this work, we aim to dig deeper into each direction in order to identify challenges and shortcomings of each study and propose unexplored avenues for addressing them.
@@
The rest of this survey is organized as follows: In Section 2, we discuss multi-modal feature study by introducing some widely spread categories of misinformation in multi-modal settings and commonly used cross-modal clues for detecting them.In the following section, we discuss different fusion mechanisms to merge modalities that are involved in such clues. Then, we explain multi-modal model study by introducing solutions and categorizing them based on the machine learning techniques they utilize. In Section 4, we describe the multi-modal data study by introducing, analyzing and comparing existing databases for multi-modal fake news detection. Finally, we discuss existing challenges and shortcomings that each direction is facing in Section 5, and propose new avenues for addressing those shortcomings and furthering the multi-modal misinformation detection research.
@@
We conducted our literature search across multiple databases, including IEEE Xplore, ACM Digital Library. and Google Scholar, using a combination of keywords related to our research focus. The inclusion criteria for the papers were defined by relevance to the research question, publication date within the last ten years to ensure timeliness, and peer-reviewed status to guarantee quality. The selection process involved an initial screening of titles and abstracts, followed by a full-text review to confirm that each paper met our stringent criteria. This methodical approach ensures that the included papers provide a diverse yet focused perspective on the subject, offering readers a succinct and informative summary of current knowledge in the field. We acknowledge the importance of transparency in our literature selection process and outline these steps to clarify the criteria and rationale behind our choices.
@@ 2. Multi-modal feature study
In this section, we discuss feature based direction of multi-modal misinformation study. To better understand the rationale behind multi-modal features and clues, we start with a brief introduction of some of the common categories of misinformation that spread in multi-modal environments. Furthermore, we discuss some of the commonly used multi-modal features and clues and then we talk about existing fusion mechanisms for combining data modality features. Finally, we discuss pros and cons of each fusion mechanism.
@@ 2.1. Common categories of misinformation in multi-modal environments
Multi-modal misinformation refers to a package of misleading information which includes multiple modalities such as image, text, video and so on and so forth. In multi-modal misinformation, not all modalities are necessarily false, but sometimes the connections between the modalities are manipulated to deceive the perception of audiences. In what follows, we briefly discuss some of the common categories of misinformation that are widely spread in multi-modal settings. It is worth mentioning that these categories of misinformation are common types of misinformation in both multi-modal and uni-modal environments. However, we provide examples of each category in multi-modal platforms as well.
@@
  • Satire or parody: This category refers to content that convey true information with a satirical tone or added information that makes it false. One of the well-known publishers of this category is TheOnion website [^2] https://www.theonion.com/ which is a digital media organization that publishes satirical articles on a variety of international, national, and local news. A multi-modal example of this category is an image within a satirical news article that contains absurd or ridiculous content or is manipulated to create humorous critique (Li et al., 2020b; Das, 2019). In this case, the textual content may not necessarily be false, but when combined with an image, it creates misleading content.
@@
  • Fabricated content: This is a category of information which is completely false and is generated to deceive the audience. The intention to publish fabricated content is usually to mislead people for political, social or economic benefits. A multi-modal instance of this category is a news report that uses auxiliary images or videos that are either completely fake or belong to irrelevant events.
@@
  • Imposter content: This is a category of misinformation that takes advantage of established news agencies by publishing misleading content under their branding. Since audiences trust established agencies, they are less likely to doubt the validity of content and consequently pay less attention to subtle clues. Imposter content may damage the reputation of agencies and undermines audiences’ trust. An example of imposter content is a website that mimics the domain features of global news outlets, such as CNN and BBC. To detect this category of misinformation, it is very crucial to identify and pay attention to subtle features of web publishers (Abdali et al., 2020, 2021a).
@@
  • Manipulated content: This is a category of misinformation that is generated by editing a valid information usually in form of images and videos to deceive the audiences. Deepfake videos are well-known examples of this category. Manipulated videos and images have been widely generated to support fabricated contents (Tolosana et al., 2020; Rössler et al., 2019).
@@
  • False connection: This is one of the most common types of misinformation in multi-modal environments. In this category, some of the modalities such as captions or titles do not support other modalities e.g., text or video. False connection is designed to catch the audience’s attention with clickbait headlines or provocative images (Nakamura et al., 2020; Li et al., 2020a).
@@ 2.2. Multi-modal features and clues
As previously indicated, combining features, such as text and image, has recently been utilized to identify false information in multi-modal contexts. In this section, we provide a non-exhaustive list of frequently used cues that machine learning researchers have used to identify false information. We emphasize that even though there are numerous other multi-modal combinations, they have not yet been fully explored by researchers at the time of writing and we merely enumerate those that are frequently used in the literature.
@@
Combination of textual content and article image is one of the widely used sets of features for multi-modal fake news detection. The intuition behind this cue is: some fake news spreaders use tempting images e.g., exaggerated, dramatic or sarcastic graphics which are far from the textual content to attract users’ attention. Since it is difficult to find both pertinent and pristine images to match these fictions, fake news generators sometimes use manipulated images to support non-factual scenarios. Researchers refer to this cue as the similarity relationship between text and image (Zhou et al., 2020b; Giachanou et al., 2020; Xue et al., 2021) which could be captured with a variety of similarity measuring techniques such as cosine similarity between the title and image tags embeddings (Zhou et al., 2020b; Giachanou et al., 2020) or similarity measure architectures (Xue et al., 2021).
@@
In video-based platforms such as YouTube, TikTok, etc., the video content is served with descriptive textual information such as video description, title, users’ comments, and replies. Different users and video producers use different writing styles in such textual content. These writing styles can be learned and distinguished from unrecognized patterns by machine learning models. Meanwhile, the meaningful relationship between the visual content and the descriptive information, e.g., the video title, is another important clue that could be used for detecting online misbehavior  (Choi and Ko, 2022). However, this is a very challenging task, as it is difficult to detect frames that are relevant to the text and discard irrelevant ones, e.g., advertisements, opening, or ending frames. Moreover, encoding all video frames is very inefficient in terms of speed and memory.
@@
The majority of the online fact checkers such as BS Detector [^6] https://github.com/selfagency/bs-detector or News Guard [^7] https://www.newsguardtech.com/ provide labels that pertain to domains rather than articles. Despite this disparity, there are several works (Helmstetter and Paulheim, 2018; Zhou, 2017) that show the weakly-supervised task of using labels pertaining to domains, and subsequently testing on labels pertaining to articles, yields negligible accuracy loss due to the strong correlation between the two (Helmstetter and Paulheim, 2018; Zhou, 2017). Thus, by recognizing the domain features and behaviors, we might be able to classify articles published by them with admissible accuracy. Some of these feature patterns are the propagation network and word usage patterns of the domains which could be considered (Silva et al., 2021b; Shu et al., 2019b; Silva et al., 2021a; Zhou and Zafarani, 2019) as a discriminating signature for different domains. It has been empirically shown that not only do news articles from different domains have significantly different word usage, but also they follow different propagation patterns (Silva et al., 2021b).
@@
Another domain level feature that researchers have recently introduced for detecting misinformation is the overall look of serving webpage (Abdali et al., 2020, 2021a). It is shown that, in contrast to credible domains, unreliable web-based news outlets tend to be visually busy and full of events such as advertisements, popups etc (Abdali et al., 2021a). Trustworthy webpages often look professional and ordered, as they often request users to agree to sign up or subscribe, have some featured articles, a headline picture, standard writing styles and so on. On the other hand, unreliable domains tend to have unprofessional blog-post style, negative space and sometimes hard-to-read font errors. Considering this discriminating clue, researchers have recently proposed to consider the overall look of the webpages in addition to textual content and social context in order to create a multi-modal model for detecting misinformation (Abdali et al., 2020, 2021b).
@@
Due to the ubiquity of camera devices and video-editing applications, video-based frameworks are extremely vulnerable to manipulation, e.g., virtual backgrounds, anime filters, etc. Such visual manipulations introduce a non-trivial noise to the video frames which may lead to misclassification of irrelevant information from videos (Shang et al., 2021). Moreover, manipulated videos often incorporate content in different modalities such as audio and text which sometimes none of them is misinformative while considered individually. However, they mislead the audiences while considering jointly with the video content. In order to detect misleading content that is jointly expressed in video, audio, and text content, researchers have proposed to leverage frame-based information along with audio and text content on video-based platforms like TikTok (Shang et al., 2021).

### v7

@@ Abstract.
As social media platforms evolve from text-based forums into multi-modal environments, the nature of misinformation in social media is also transforming accordingly. Taking advantage of the fact that visual modalities such as images and videos are more favorable and attractive to users, and textual content is sometimes skimmed carelessly, misinformation spreaders have recently targeted contextual connections between the modalities, e.g., text and image. Hence, many researchers have developed automatic techniques for detecting possible cross-modal discordance in web-based content. We analyze, categorize, and identify existing approaches in addition to the challenges and shortcomings they face in order to unearth new research opportunities in the field of multi-modal misinformation detection.
@@ 1. Introduction
Nowadays, billions of multi-modal posts containing texts, images, videos, soundtracks, etc., are shared throughout the web, mainly via social media platforms such as Facebook, Twitter, Snapchat, Reddit, Instagram, YouTube, and so on. While the combination of modalities allows for more expressive, detailed, and user-friendly content, it brings about new challenges, as it is harder to accommodate uni-modal solutions to multi-modal environments.
@@
Similar to other multi-modal tasks, it is harder and more challenging to detect fake news on multi-modal platforms, as it requires not only the evaluation of each modality, but also cross-modal connections and credibility of the combination as well. This becomes even more challenging when each modality e.g., text or image is credible but the combination creates misinformative content. For instance, a COVID-19 anti-vaccination misinformation111“Misinformation” is false information that spreads unintentionally, whereas the term “Disinformation” refers to false information that malicious users share intentionally and often strategically to affect other audiences’ behaviors toward social, political, and economic events. In this work, regardless of spreaders’ intention, we refer to all sorts of false news i.e., misinformation and disinformation as “Misinformation” or “Fake News” interchangeably. post can have text that reads “vaccines do this” and then attaches a graphic image of a dead person. In this case, although the image and text are not individually misinformative, taken together they create misinformation.
@@
Over the past decade, several detection models (Shu et al., 2017, 2020c; Islam et al., 2020; Cai et al., 2020) have been developed to detect misinformation. However, the majority of them leverage only a single modality for misinformation detection, e.g., text (Horne and Adali, 2017; Wu et al., 2017; Guacho et al., 2018; Shu et al., 2020a) or image (Huh et al., 2018; Qi et al., 2019; Abdali et al., 2021a; Choudhary and Arora, 2021), which miss the important information conveyed by other modalities. There are existing works (K. Shu and Liu, 2019; Shu et al., 2019a; Abdali et al., 2021b, 2020; Hakak et al., 2021) that leverage ensemble methods which create multiple models for each modality and then combine them to produce improved results. However, in many cases of multi-modal misinformation, loosely combining individual modalities is inadequate for detecting fake news, leading to the failure of the joint model.
@@
Nevertheless, in recent years, machine learning scientists have developed different techniques for cross-modal fake news detection, which combine information from multiple modalities, leveraging cross-modal information such as the consistency and meaningful relationships between different modalities. Studying and analyzing these techniques and identifying existing challenges will give a clearer picture of the state of knowledge on multi-modal misinformation detection and open the door to new opportunities in this field.
@@
Even though there are a number of valuable surveys on fake news detection (Shu et al., 2017; Kumar and Shah, 2018; da Silva et al., 2019), very few of them focus on multi-modal techniques (Alam et al., 2021; Segura-Bedmar and Alonso-Bartolome, 2022). Since the number of proposed techniques for multi-modal fake news detection has been increasing immensely, the necessity of a comprehensive survey on existing techniques, datasets, and emerging challenges is felt more than ever. With that said, in this work, we aim to conduct a comprehensive study on fake news detection in multi-modal environments.
@@
To this end, we classify multi-modal misinformation detection study into following directions:
@@
  • Multi-modal Data Study: in this direction, the goal is to collect multi-modal fake news data, e.g., image, text, social context, etc., from different sources of information and use fact-checking resources to evaluate the veracity of the collected data and annotate them accordingly. Comparison and analysis of existing datasets, as well as benchmarking, are other tasks that fall under this category.
@@
  • Multi-modal Feature Study: the primary goal of this study is to uncover significant links between various data modalities, which are frequently exploited by misinformation spreaders to distort, impersonate, or exaggerate original information. These meaningful connections may be used as clues for detecting misinformation in multi-modal environments such as social media posts. Another goal of this direction is to study and develop strategies for fusing features of different modalities and creating information-rich multi-modal features.
@@
  • Multi-modal Model Study: the main focus of this direction is on the development of efficient multi-modal machine learning solutions to detect misinformation by leveraging multi-modal features and clues. Proposing new techniques and approaches, in addition to improving the performance, scalability, interpretability, and explicability of machine learning models, are some of the common tasks in this direction.
@@
These three studies form a sequential pipeline in the multi-modal misinformation field, where the output of each study serves as the input for the next. Fig. 1 provides a summary of these directions. In this work, we aim to explore each direction in greater depth to identify the challenges and shortcomings of each study and propose new avenues for addressing them.
@@
The rest of this survey is organized as follows: In Section 2, we discuss the multi-modal feature study by introducing some widely spread categories of misinformation in multi-modal settings and commonly used cross-modal clues for detecting them. In the following section, we discuss different fusion mechanisms to merge modalities involved in such clues. Then, we explain the multi-modal model study by introducing solutions and categorizing them based on the machine learning techniques they utilize. In Section 4, we describe the multi-modal data study by introducing, analyzing, and comparing existing databases for multi-modal fake news detection. In Section 5, we discuss existing challenges and shortcomings that each direction is facing. Finally, in Section 6, we propose new avenues to address these shortcomings and advance multi-modal misinformation detection research.
@@
We conducted our literature search across multiple databases, including IEEE Xplore, ACM Digital Library, and Google Scholar, using a combination of keywords related to our research focus. The inclusion criteria for the papers were defined by their relevance to the research question, publication date within the last ten years to ensure timeliness, and peer-reviewed status to guarantee quality. The selection process involved an initial screening of titles and abstracts, followed by a full-text review to confirm that each paper met our stringent criteria. This methodical approach ensures that the included papers provide a diverse yet focused perspective on the subject, offering readers a succinct and informative summary of current knowledge in the field. We emphasize the importance of transparency in our literature selection process and outline these steps to clarify the criteria and rationale behind our choices.
@@ 2. Multi-modal feature study
In this section, we discuss the feature-based direction of multi-modal misinformation studies. To better understand the rationale behind multi-modal features and clues, we start with a brief introduction to some of the common categories of misinformation that spread in multi-modal environments. Furthermore, we discuss some of the commonly used multi-modal features and clues, and then we talk about existing fusion mechanisms for combining data modality features. Finally, we discuss the pros and cons of each fusion mechanism.
@@ 2.1. Common Categories of Misinformation in Multi-modal Environments
Multi-modal misinformation refers to a package of misleading information that includes multiple modalities such as images, text, videos, and so on. In multi-modal misinformation, not all modalities are necessarily false, but sometimes the connections between the modalities are manipulated to deceive the audience’s perception. In what follows, we briefly discuss some of the common categories of misinformation that are widely spread in multi-modal settings. It is worth mentioning that these categories of misinformation are common in both multi-modal and uni-modal environments. However, we provide examples of each category in multi-modal platforms as well.
@@
  • Satire or Parody: This category refers to content that conveys true information with a satirical tone or added information that makes it false. One of the well-known publishers of this category is The Onion website [^2] https://www.theonion.com/, which is a digital media organization that publishes satirical articles on a variety of international, national, and local news. A multi-modal example of this category is an image within a satirical news article that contains absurd or ridiculous content or is manipulated to create humorous critique (Li et al., 2020b; Das, 2019). In this case, the textual content may not necessarily be false, but when combined with an image, it creates misleading content.
@@
  • Fabricated Content: This category of information is completely false and is generated to deceive the audience. The intention behind publishing fabricated content is usually to mislead people for political, social, or economic benefits. A multi-modal instance of this category is a news report that uses auxiliary images or videos that are either completely fake or belong to irrelevant events.
@@
  • Imposter Content: This category of misinformation takes advantage of established news agencies by publishing misleading content under their branding. Since audiences trust established agencies, they are less likely to doubt the validity of the content and consequently pay less attention to subtle clues. Imposter content may damage the reputation of agencies and undermine audience trust. An example of imposter content is a website that mimics the domain features of global news outlets, such as CNN  [^3] https://www.cnn.com/world and BBC  [^4] https://www.bbc.com/. To detect this category of misinformation, it is crucial to identify and pay attention to the subtle features of web publishers (Abdali et al., 2020, 2021a).
@@
  • Manipulated Content: This category of misinformation is generated by editing valid information, usually in the form of images and videos, to deceive audiences. Deepfake videos are well-known examples of this category. Manipulated videos and images have been widely generated to support fabricated content (Tolosana et al., 2020; Rössler et al., 2019; Abdali et al., 2021c).
@@
  • False Connection: This is one of the most common types of misinformation in multi-modal environments. In this category, some modalities, such as captions or titles, do not support other modalities, such as text or video. False connections are designed to catch the audience’s attention with clickbait headlines or provocative images (Nakamura et al., 2020; Li et al., 2020a).
@@ 2.2. Multi-modal Features and Clues
As previously indicated, combining features such as text and images has recently been utilized to identify false information in multi-modal contexts. In this section, we provide a non-exhaustive list of frequently used cues that machine learning researchers have used to identify false information. We emphasize that even though there are numerous other multi-modal combinations, they have not yet been fully explored by researchers at the time of writing, and we merely enumerate those that are frequently used in the literature.
@@
The combination of textual content and article images is one of the widely used sets of features for multi-modal fake news detection. The intuition behind this cue is that some fake news spreaders use tempting images, such as exaggerated, dramatic, or sarcastic graphics, which are far from the textual content to attract users’ attention. Since it is difficult to find both pertinent and pristine images to match these fictions, fake news generators sometimes use manipulated images to support non-factual scenarios. Researchers refer to this cue as the similarity relationship between text and image (Zhou et al., 2020b; Giachanou et al., 2020; Xue et al., 2021), which could be captured with a variety of similarity measuring techniques such as cosine similarity between the title and image tags embeddings (Zhou et al., 2020b; Giachanou et al., 2020) or similarity measure architectures (Xue et al., 2021).
@@
On video-based platforms such as YouTube  [^8] https://www.youtube.com/ and TikTok  [^9] https://www.tiktok.com/, video content is accompanied by descriptive textual information such as video descriptions, titles, users’ comments, and replies. Different users and video producers use various writing styles in such textual content. These writing styles can be learned and distinguished from unrecognized patterns by machine learning models. Meanwhile, the meaningful relationship between the visual content and the descriptive information, such as the video title, is another important clue that could be used for detecting online misbehavior (Choi and Ko, 2022). However, this is a very challenging task, as it is difficult to detect frames that are relevant to the text and discard irrelevant ones, such as advertisements, opening, or ending frames. Moreover, encoding all video frames is very inefficient in terms of speed and memory.
@@
The majority of online fact checkers, such as BS Detector [^10] https://github.com/selfagency/bs-detector or News Guard [^11] https://www.newsguardtech.com/, provide labels that pertain to domains rather than articles. Despite this disparity, several works (Helmstetter and Paulheim, 2018; Zhou, 2017) show that the weakly-supervised task of using labels pertaining to domains and subsequently testing on labels pertaining to articles yields negligible accuracy loss due to the strong correlation between the two (Helmstetter and Paulheim, 2018; Zhou, 2017). Thus, by recognizing the domain features and behaviors, we might be able to classify articles published by them with admissible accuracy. Some of these feature patterns are the propagation network and word usage patterns of the domains, which could be considered (Silva et al., 2021b; Shu et al., 2020b; Silva et al., 2021a; Zhou and Zafarani, 2019) as a discriminating signature for different domains. It has been empirically shown that not only do news articles from different domains have significantly different word usage, but they also follow different propagation patterns (Silva et al., 2021b).
@@
Another domain-level feature that researchers have recently introduced for detecting misinformation is the overall look of the serving webpage (Abdali et al., 2020, 2021a). It is shown that, in contrast to credible domains, unreliable web-based news outlets tend to be visually busy and full of events such as advertisements, popups, etc. (Abdali et al., 2021a). Trustworthy webpages often look professional and ordered, as they often request users to agree to sign up or subscribe, have some featured articles, a headline picture, standard writing styles, and so on. On the other hand, unreliable domains tend to have an unprofessional blog-post style, negative space, and sometimes hard-to-read font errors. Considering this discriminating clue, researchers have recently proposed to consider the overall look of the webpages in addition to textual content and social context in order to create a multi-modal model for detecting misinformation (Abdali et al., 2020, 2021b).
@@
Due to the ubiquity of camera devices and video-editing applications, video-based frameworks are extremely vulnerable to manipulation, e.g., virtual backgrounds, anime filters, etc. Such visual manipulations introduce non-trivial noise to the video frames, which may lead to the misclassification of irrelevant information from videos (Shang et al., 2021). Moreover, manipulated videos often incorporate content in different modalities such as audio and text, which sometimes are not misinformative when considered individually. However, they mislead the audience when considered jointly with the video content. To detect misleading content that is jointly expressed in video, audio, and text content, researchers have proposed leveraging frame-based information along with audio and text content on video-based platforms like TikTok (Shang et al., 2021).